Calculate the wavelength of the following “particles”:

(a) A 6.00 x 10-2 kg tennis ball traveling at 68 m/s

(b) An electron moving at the same speed


1
Expert's answer
2022-01-12T09:26:02-0500

Use De- - broglie's wavelength formula ,

λ = h/P , here h is plank's constant and P is linear momentum of particle


Given, mass of body , m = 6 × 10⁻² Kg

velocity of body , v = 62 m/s

so, momentum , P = mass × velocity = 6 × 52 × 10⁻² Kgm/s

= 3.12 Kgm/s


Now, wavelength , λ = 6.626 × 10⁻³⁴/3.12 m [ plank's constant, h = 6.626 × 10⁻³⁴Js]

λ = 2.1 × 10⁻³⁴ m , here you can see wavelength associated by tennis ball is so small , negligible . So we should consider particle nature . That's what said de - broglie's in his Theory , " any object have dual nature , particle as well as wave . When object is very small and speed of it is very high then it considered as a wave because wavelength will be able to compare . But when particle is large and speed of it is low then it considered as particle .